OpenStudy (anonymous):

Repeat the following procedure for the four given numbers Multiply the number by 4. add 12 to the product. divide this sum by 2. subtract 6 from the quotient 1st number is 3 2nd number is 6 3rd number is 8 4th number is 12 Write a conjecture that relates the result of the process to the original number selected. represent the original number as N the result is?

OpenStudy (anonymous):

call then number n first step \[4n\] second step \[4n+12\] third step \[\frac{4n+12}{2}=2n+6\] fourth step \[2n+6-6=2n\]

OpenStudy (anonymous):

SO this is a long way to multiply a number by 2.